Determine what the next equation would be, and verify that it is indeed a true statement.
6)

(1 × 9) - 4 = 5
(21 × 9) - 4 = 185
(321 × 9) - 4 = 2885


Answer by stanbon(75887)

Determine what the next equation would be, and verify that it is indeed a true statement.
6)
(1 × 9) - 4 = 5
(21 × 9) - 4 = 185
(321 × 9) - 4 = 2885
-------------------
(4321 x 9)-4 = 38885
============================
Cheers,
Stan H.
=======